OpenStudy (anonymous):

The average of three numbers is 94. The average of four different numbers is 87. What is the weighted mean of all the numbers together?

OpenStudy (anonymous):

total of first three nos =94*3=282 total of four different numbers =87*4=348 thus the total of seven numbers =282+348=630 thus the reqd mean is 630/7=90